OpenStudy (anonymous):

how do you solve log base 64 of 0.5 = x?

OpenStudy (anonymous):

definition of log is base ^ x = 0.5, so 64^x = 0.5 take the log of both sides: log(64^x) = log(0.5) . you can pull exponents in logs into the front: x log(64) = log(0.5) x = log(0.5)/log(64) x= -1/6 Check 64^(-1/6) = 1/2 :) Hope that helps!

OpenStudy (anonymous):

i don't understand this :/

OpenStudy (anonymous):

if i have log base x = a, that means base ^a = x Does that much make sense? If so I can keep going :)

OpenStudy (anonymous):

ok yea

OpenStudy (anonymous):

okay, so plugging your numbers, you have 64 ^x = 0.5 We are interested in solving for x, and there is a tricky way of doing that.

OpenStudy (anonymous):

Notice that at the moment, the x is part of an expression that is kind of weird to undo since it is the exponent. To undo it, we use a function that allows us to drag down the exponnet as a coefficient of the expression. So as review, if I have: log(a^b), this is the same as b *log(a) Does that much make nse?

OpenStudy (anonymous):

ok yeaaaaaa

OpenStudy (anonymous):

okay yay. so now we apply that idea to what we were doing. If 64 ^x = (0.5) I can take the log of both sides log(64^x) = log(0.5) And similar to the general case I explained, this is x * log(64) = log(0.5) Yeah?:)

OpenStudy (anonymous):

ok yesssssss

OpenStudy (anonymous):

wow my computer just froze, let me type that again

OpenStudy (anonymous):

so now we are just solving for x, which is totally possible because the x term is now isolated. we can think of log (64) and log (0.5) like numbers because the log of a number is another number. So we can divide each side by log(64) x*log(64) = log(0.5) x = log(0.5)/log(64) And then no one knows those things off the top of their head unless they work with logs a lot, so the typical person will evaluate the right hand side expression using a calculator x = -1/6 I hope that makes sense now!

OpenStudy (anonymous):

okay i get it now ! thank you so much !(:

OpenStudy (anonymous):

:) No problem!